For which of the following salts is there a positive entropy change for dissolution in water?

NH4NO3 SrCO3
S⁰ (s) (J/mol⋅K) 151.1 97.1
S⁰ (aq) (J/mol⋅K) 259.8 -85.7

Group of answer choices

Neither NH4NO3 nor SrCO3

Both NH4NO3 and SrCO3

NH4NO3

SrCO3
Answer #1

NH4NO3:

entropy change = S⁰ (aq) - S⁰ (s) = (259.8 - 151.1) = 108.7 J/mol-K (positive)

SrCO3:

entropy change = S⁰ (aq) - S⁰ (s) = (- 85.7 - 97.1) = - 182.8 J / mole-K (negative)

answer is NH4NO3